OpenStudy (anonymous):

a. If AB = 5y - 4, BD = 7y + 3 and AD = 20y - 9, find the value of y.

OpenStudy (einstein):

AB = 5y - 4 BD = 7y + 3 and AD = 20y - 9 AD = AB + BD 20y - 9 = 5y - 4 + 7y + 3 8y = 8 y =1